# OffLight: An Offline Multi-Agent Reinforcement Learning Framework for Traffic Signal Control

### v1

@@ Abstract
Efficient traffic signal control (TSC) is essential for modern urban mobility, but traditional systems often struggle to adapt to the complex and dynamic nature of city traffic. While Multi-Agent Reinforcement Learning (MARL) offers promising adaptive solutions, online MARL methods require a significant amount of interactions with the environment, which can be expensive and time consuming. Offline MARL addresses these concerns by leveraging historical traffic data for training, but it faces challenges due to the heterogeneity of behavior policies in real-world datasets—a mix of different controllers makes learning difficult. We introduce OffLight, a novel offline MARL framework specifically designed to handle heterogeneous behavior policies within TSC datasets. OffLight employs Gaussian Mixture Model Variational Graph Autoencoder (GMM-VGAEs) to model the complex distribution of behavior policies, enabling effective learning from diverse data sources. To enhance coordination between agents, we integrate Graph Attention Networks (GATs), allowing agents to make informed decisions based on aggregated information from neighboring intersections. Furthermore, OffLight incorporates Importance Sampling (IS) to correct for differences between the behavior and target policies and utilizes Return-Based Prioritized Sampling (RBPS) to focus on high-quality experiences, thereby improving sample efficiency. Extensive experiments across three real-world urban traffic scenarios—Jinan (12 intersections), Hangzhou (16 intersections), and Manhattan (196 intersections)—demonstrate that OffLight significantly outperforms existing offline RL methods. Notably, OffLight achieves up to a 7.8% reduction in average travel time and an 11.2% decrease in queue length compared to baseline algorithms, particularly in datasets with mixed-quality data. Our ablation studies confirm the effectiveness of OffLight’s components in handling data heterogeneity and enhancing learning performance. These results highlight OffLight’s ability to accurately model heterogeneous behavior policies, mitigate the impact of suboptimal data, and scale to large urban networks. By addressing the critical challenges of offline MARL in TSC, OffLight offers a practical and impactful solution for improving urban traffic management without the risks associated with online learning.
Keywords: Multi-Agent Reinforcement Learning, Offline Reinforcement Learning, Traffic Signal Control, Intelligent Transportation Systems, Deep Reinforcement Learning.
1 Introduction
Efficient traffic signal control (TSC) is important for modern urban mobility, directly affecting congestion levels, travel times, and overall city livability. As urban populations grow and vehicular usage intensifies, traditional traffic management systems struggle to adapt to dynamic and complex traffic flows. Recent advancements in Multi-Agent Reinforcement Learning (MARL) offer promising solutions by enabling decentralized, adaptive, and intelligent control of traffic signals [1, 2].
@@ 1 Introduction
While online MARL requires agents to interact with the environment in real-time—which can be impractical and unsafe in real-world traffic systems—offline MARL leverages historically collected traffic data to train agents without live experimentation [3]. This approach offers several advantages:
  • Extensive Data Utilization: Utilizes pre-collected traffic data, allowing agents to learn from diverse scenarios and rare events difficult to replicate online [4].
  • Safety and Risk Mitigation: Eliminates risks associated with deploying untested policies in live traffic, ensuring only well-trained agents are implemented [5].
  • Cost Efficiency and Scalability: Reduces the need for expensive simulations or field tests, facilitating deployment across various scales and environments [3].
  • Accelerated Policy Development: Enables rapid iteration and refinement of control policies, expediting research and development [6].
Several offline reinforcement learning (RL) methods have been explored for TSC, often incorporating imitation learning to bootstrap training. For instance, DemoLight uses demonstrations from traditional controllers to initialize policies, enabling efficient exploration and faster convergence compared to pure RL approaches [7]. Cooperative Control for multi-intersection TSC combines imitation learning and deep RL to coordinate multiple intersections, highlighting the benefits of learning from expert demonstrations [8]. CrossLight begins with offline training using behavior cloning and transitions to online RL to adapt policies to new environments, showcasing the effectiveness of imitation learning for strong policy initialization [9]. Building upon these efforts, DataLight introduces an input space modeling technique that captures the spatial distribution of traffic using self-attention mechanisms, demonstrating improved policy performance without requiring live interactions [4]. Figure 1 represents a general framework of an offline RL algorithm.
However, a key challenge in offline MARL for TSC remains unaddressed: real-world datasets often consist of a heterogeneous mix of behavior policies, making it difficult for agents to learn effective strategies. This heterogeneity can arise from a variety of factors, such as:
  • Local Heuristics: Many traffic signals operate on traditional rule-based controllers (e.g., fixed-time or reactive control). These heuristics are often suboptimal and vary between intersections [10].
  • Varying Control Strategies: Over time, traffic systems may switch between different controllers (e.g., from manual control to AI-based systems), resulting in datasets that reflect a blend of expert and suboptimal policies [11].
  • Temporal Variability: Traffic patterns fluctuate throughout the day, leading to variations in the policies being implemented during different time periods. For instance, a traffic signal might switch between morning peak-time strategies and off-peak strategies [10].
This heterogeneity hampers effective learning because:
  • Offline MARL methods are highly sensitive to the underlying distribution of policies in the dataset. Mixing high- and low-quality data complicates the learning process, leading to poor generalization [12].
  • Mixed-policy data can introduce biases, making it harder for the agent to identify the most effective policies [13].
To address these challenges, we introduce OffLight, a novel offline MARL framework that combines Importance Sampling (IS) and Return-Based Prioritized Sampling (RBPS) to mitigate distributional shifts and focus on high-value experiences. This synergy between IS and RBPS enhances both sample efficiency and policy performance.
GMM-VGAE models the diversity in behavior policies within the traffic network, capturing policy heterogeneity by disentangling them in the latent space. By leveraging Graph Neural Networks (GNNs), OffLight combines local observations across intersections into a structured global representation, which is crucial in traffic signal control where agent interactions are highly localized but collectively influence network-wide patterns. This enables OffLight to use IS effectively, correcting for distributional shifts and ensuring stable policy learning from varied data. Meanwhile, RBPS prioritizes high-return episodes, allowing OffLight to focus on the most informative experiences, thus accelerating convergence and optimizing performance.
Key contributions of this paper include:
  • Developing OffLight, an offline MARL framework designed to handle heterogeneous behavior policies in traffic signal control (TSC) using a combination of Importance Sampling (IS) and Return-Based Prioritized Sampling (RBPS).
  • Introducing GMM-VGAE to accurately capture diverse behavior policies, supporting robust learning under policy heterogeneity and enabling effective IS.
  • Leveraging RBPS to prioritize high-return episodes, which enhances sample efficiency and accelerates convergence by focusing on the most informative experiences.
  • Demonstrating OffLight’s scalability and superior performance on real-world traffic datasets, where it consistently outperforms existing methods.
The remainder of the paper is organized as follows. Section 2 reviews related work and provides necessary background on MARL and offline RL in the context of TSC. Section 3 details the OffLight framework, including its architecture and key components. Section 4 describes the experimental setup, and Section 5 presents the results and analysis. Finally, Section 6 concludes the paper and Section LABEL:sec:discussion discusses the key insights gained from this study and limitations of OffLight.
2 Background
2.1 Multi-Agent Reinforcement Learning
Multi-Agent Reinforcement Learning (MARL) involves multiple agents interacting within a shared environment. Each agent learns to maximize the cumulative reward while adapting to the behaviors of other agents. MARL introduces additional challenges such as non-stationarity (due to concurrent learning), the need for coordination, and partial observability, all of which make policy learning significantly harder than in single-agent settings [14, 15].
2.1.1 Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs)
Multi-agent problems are commonly modeled as Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs) [16]. A Dec-POMDP is formally defined by a tuple $\langle N,\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{N},\{\mathcal{O}^{i}\}_{i=1}^{N},P,R,H,\gamma\rangle$, where:
  • $N$ is the number of agents.
  • $\mathcal{S}$ is the set of environment states.
  • $\mathcal{A}^{i}$ is the action space of agent $i$.
  • $\mathcal{O}^{i}$ is the observation space of agent $i$.
  • $P:\mathcal{S}\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{N}\times\mathcal{S}\rightarrow[0,1]$ is the state transition probability function.
  • $R:\mathcal{S}\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{N}\rightarrow\mathbb{R}$ is the shared reward function.
  • $H$ is the finite horizon.
  • $\gamma\in[0,1)$ is the discount factor.
Each agent $i$ selects actions based on its action-observation history $h^{i}_{t}=\{o^{i}_{0},a^{i}_{0},\dots,o^{i}_{t-1},a^{i}_{t-1},o^{i}_{t}\}$, where $o^{i}_{t}\in\mathcal{O}^{i}$ is the observation at time $t$, and $a^{i}_{t}\in\mathcal{A}^{i}$ is the action. The joint action $\mathbf{a}_{t}=(a_{t}^{1},\dots,a_{t}^{N})$ determines the next state $s_{t+1}$ according to $P$, and all agents receive a shared reward $R$. The objective in Dec-POMDPs is to find a set of decentralized policies $\{\pi^{i}\}_{i=1}^{N}$, where each $\pi^{i}:H^{i}_{t}\rightarrow\mathcal{A}^{i}$ maps action-observation histories to actions, such that the expected cumulative reward is maximized:
@@ 2.1.1 Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs)
2.2 Application to Traffic Signal Control (TSC)
In the context of Traffic Signal Control (TSC), each traffic signal controller can be modeled as an agent within the Decentralized Partially Observable Markov Decision Process (Dec-POMDP) framework [16]. Each agent operates independently and makes decisions based on local observations, such as vehicle counts, vehicle speeds, queue lengths, and current signal phases.
Observation Space: Each traffic signal agent has a limited range of observation. This range reflects the realistic coverage of common sensors such as inductive loop detectors or cameras. For each incoming lane, the agent observes:
  • The number of vehicles $\{n_{l}\}_{l=1}^{L_{i}}$, where $L_{i}$ is the number of lanes for traffic signal $i$.
  • The average speed of vehicles $\{s_{l}\}_{l=1}^{L_{i}}$, normalized by the speed limit.
  • The number of halted vehicles, representing the queue length $\{q_{l}\}_{l=1}^{L_{i}}$.
  • The current phase ID, indicating the traffic signal’s current phase.
These local observations provide the sensory input required for each agent to make real-time decisions about the traffic signal phases.
Action Space: Each traffic signal agent controls the signal phases at its intersection. The available actions correspond to selecting one of the predefined green phases from the list of possible phases for that intersection. An agent may select a green phase and it must then transition to the yellow phase, which is enforced by the environment. The action selection interval is fixed at 5 simulation seconds.
Reward Function: A commonly used reward metric is queue length. Minimizing the queue length across all lanes in the network directly reflects the reduction of vehicle stoppage. The global reward at each time step $r_{t}$ is defined as:
$r_{t}=-\sum_{l\in L}q_{l}(t),$
where $q_{l}(t)$ represents the queue length on lane $l$ at time $t$, and $L$ is the set of all lanes in the network. This reward structure encourages agents to minimize congestion and improve traffic flow throughout the system.
Deep MARL: In multi-agent TSC, the actions of one traffic signal directly impact the observations and performance of neighboring intersections. Therefore, coordination between agents is crucial to achieve optimal traffic management [17]. Traditional approaches, such as Independent Q-Learning (IQL) [18], treat each agent as independent, which may lead to suboptimal coordination. Recent advancements in the application of deep MARL in TSC include frameworks such as CoLight [11], PressLight [19], AttendLight [20] which utilize Graph Attention Networks (GATs) [21, 11] to model inter-agent communication. These frameworks model the traffic network as a graph, where intersections are treated as nodes and roads as edges, allowing traffic signals to share information with neighboring intersections. By focusing attention on the most relevant neighbors, these methods have achieved significant success in optimizing traffic flow through coordinated agent behavior.
Following the success of these frameworks in the online MARL setting, we choose this approach as the foundation for our framework to model offline MARL policies. By leveraging the well-established GAT-based communication mechanism, we aim to maintain effective coordination between agents, even when learning occurs from static offline datasets. This approach ensures that OffLight can efficiently handle the complexity of real-world traffic systems, while addressing the challenges inherent in offline learning.
2.3 Offline Multi-Agent Reinforcement Learning
Offline Multi-Agent Reinforcement Learning involves learning effective policies from a fixed dataset of interactions without further environmental exploration during training [3, 4]. This setting is crucial in scenarios where data collection is expensive, risky, or impractical, such as traffic signal control. Offline MARL introduces unique challenges:
  • Distributional Shift: The static dataset may not adequately represent the state-action distribution encountered when executing the learned policies, leading to overestimation of out-of-distribution actions and degraded performance [22].
  • Extrapolation Error: Without any interaction with the environment, agents must generalize based solely on the dataset provided. This limitation risks extrapolation errors, where agents might attempt actions unsupported by the data, leading to poor or unsafe decisions when deployed.
  • Limited Coverage of Optimal Strategies: Offline datasets may lack the comprehensive coverage needed for optimal joint policies. This is particularly challenging in MARL, as optimal coordination patterns often require data from multiple interacting policies, which may be missing in historical data.
Recent offline RL algorithms, such as Conservative Q-Learning (CQL) [12] and TD3+BC [23], have been developed to address these challenges by regularizing learned policies and penalizing overestimation of Q-values for unseen actions. While originally designed for single-agent settings, these methods can be extended to multi-agent systems using frameworks like Centralized Training with Decentralized Execution (CTDE) [24].
2.3.1 Conservative Q-Learning (CQL)
Conservative Q-Learning (CQL) [12] is an offline RL algorithm designed to address the overestimation of out-of-distribution (OOD) actions by learning conservative Q-value estimates. CQL modifies the standard Bellman error minimization by adding a regularization term that penalizes Q-values for actions not present in the dataset. The CQL objective is: {dmath} min_Q α( E_s ∼D, a ∼π(a|s) [Q(s,a)] - E_(s,a) ∼D [Q(s,a)] ) + L_Bellman(Q), where $\mathcal{D}$ is the offline dataset, $\pi(a|s)$ is a policy (e.g., the uniform policy), $\alpha$ is a regularization coefficient, and $\mathcal{L}_{\text{Bellman}}(Q)$ is the standard Bellman error. In multi-agent settings, the joint action space grows exponentially with the number of agents, making it infeasible to represent the full joint Q-function. To address this, CQL can be extended using factorization methods or the CTDE paradigm [24]. CTDE allows agents to be trained centrally with access to global information, while at execution time, they act based on local observations, enabling scalability and coordination among agents even with limited data.
2.3.2 TD3+BC
TD3+BC [23] enhances the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm [25] by integrating a behavior cloning (BC) term into the policy update. This addition encourages the learned policy to stay close to the behavior policy present in the offline dataset, mitigating issues arising from distributional shift. The modified policy objective in TD3+BC is:
$\min_{\pi_{\theta}}-\mathbb{E}_{s\sim\mathcal{D}}[Q_{\phi}(s,\pi_{\theta}(s))]+\lambda\mathbb{E}_{(s,a)\sim\mathcal{D}}[\|\pi_{\theta}(s)-a\|_{2}^{2}],$ (2)
where $\pi_{\theta}$ is the policy network parameterized by $\theta$, $Q_{\phi}$ is the critic network parameterized by $\phi$, and $\lambda$ balances the importance of the behavior cloning term. In multi-agent systems, TD3+BC can be adapted by allowing each agent to learn its own policy while considering the interactions with other agents. Using the CTDE framework [26], agents are trained with access to additional global information (e.g., other agents’ actions or policies) to improve coordination. During execution, agents operate based on their local observations, ensuring scalability and feasibility in decentralized environments.
2.4 Challenge of Heterogeneous Behavior Policies in Offline MARL
A unique challenge in offline MARL, particularly in domains like Traffic Signal Control (TSC), is managing heterogeneous behavior policies present in the offline dataset. In TSC, data often come from various sources, times, or operational contexts, leading to a mix of behavior policies that agents must learn from. This heterogeneity adds complexity to offline MARL in the following ways:
  • Increased Risk of Distributional Shift: When policies in the offline data vary significantly, the likelihood of a distributional mismatch between the training data and the learned policy increases. The agents are more prone to encountering out-of-distribution actions, resulting in unreliable Q-value estimates and degraded policy performance, as the diversity in behavior data can make distributional correction methods less effective [22].
  • Difficulty in Behavior Policy Estimation: For methods relying on importance sampling or policy constraints, accurately estimating a single behavior policy becomes challenging when the data reflects multiple, potentially conflicting policies. Modeling this mixture accurately is essential to effective learning but is significantly more complex in offline MARL with heterogeneous datasets [13].
Addressing the challenge of heterogeneous behavior policies is essential for effective offline MARL, especially in domains like TSC where operational conditions and policies can vary widely.
2.5 Motivation for the OffLight Framework
To tackle the specific challenge of heterogeneous behavior policies in offline MARL, we propose the OffLight framework. OffLight is designed to:
  • Model Heterogeneous Behavior Policies: Utilize advanced modeling techniques, such as Variational Graph Auto-Encoders (GMM-VGAEs), to accurately estimate and represent the diverse behavior policies present in the offline dataset.
  • Mitigate Distributional Shift: Employ importance sampling with the estimated behavior policy to correct for discrepancies between the behavior policies and the target policy, reducing overestimation and improving policy evaluation.
  • Improve Sample Efficiency: In addition to IS and GMM-VGAE, OffLight integrates Return-Based Prioritized Sampling (RBPS) based on episodic returns. RBPS prioritizes high-reward episodes, ensuring that the learning process focuses on the most informative and successful experiences. This prioritization accelerates convergence and enhances the robustness of the learned policies.
By focusing on accurately modeling and leveraging the heterogeneous behavior policies in the offline data, OffLight addresses a critical challenge in offline MARL for TSC.
3 OffLight Framework

### v3

@@ Abstract
Efficient traffic signal control (TSC) is crucial for reducing congestion and improving urban mobility. While multi-agent reinforcement learning (MARL) has shown promise in adaptive traffic management, its reliance on real-time interactions makes online training costly and impractical. Offline MARL, which learns from historical data, offers a safer and more scalable alternative but struggles with heterogeneous behavior policies in real-world datasets. We introduce OffLight, a novel offline MARL framework designed to handle policy heterogeneity in traffic signal control datasets. OffLight integrates importance sampling (IS) to correct for distributional shifts and return-based prioritized sampling (RBPS) to emphasize high-quality experiences. To model diverse behavior policies, it leverages a Gaussian mixture model variational graph autoencoder (GMM-VGAE), capturing spatial and temporal traffic dynamics while improving policy estimation. We evaluate OffLight across real-world urban traffic scenarios varying from small to large traffic signal networks. Results show that OffLight outperforms existing offline RL methods, reducing average travel time by up to 7.8% and queue length by 11.2% compared to state-of-the-art approaches. Unlike prior offline RL methods, OffLight effectively adapts to complex and mixed-policy datasets, making it more reliable for real-world deployment without risky online training. These results highlight OffLight’s potential to improve urban traffic management at scale. Our implementation is available at https://github.com/rbokade/offlight.
I INTRODUCTION
Efficient traffic signal control (TSC) is critical for urban mobility, directly affecting congestion, travel times, and city livability. Multi-agent reinforcement learning (MARL) has emerged as a promising solution by enabling decentralized, adaptive, and intelligent control of traffic signals [1, 2]. While online MARL methods rely on real-time interaction, offline MARL uses pre-collected data, offering safer, more cost-efficient, and scalable alternatives by eliminating the need for risky real-time experimentation [3, 4, 5, 6].
@@ I INTRODUCTION
Several offline RL methods have shown promise in traffic signal control by using imitation learning to bootstrap initial training. Methods such as DemoLight [7], Cooperative Control [8], and CrossLight [9] have leveraged expert demonstrations to quickly initialize and refine traffic control policies. More recently, DataLight [4] improved policy performance through attention mechanisms that capture spatial traffic distributions. However, these approaches do not fully address a significant issue in offline MARL datasets: heterogeneous behavior policies arising from varied control strategies, temporal variability, and the mixture of different operational contexts [10, 11].
The heterogeneity in real-world traffic datasets poses significant challenges for offline MARL, primarily due to distributional shifts and suboptimal behavior biases that degrade policy learning [12, 13]. Recent studies, including Hong et al.[14] and Yue et al.[15], have proposed prioritized resampling techniques to emphasize high-quality experiences, thus improving policy robustness in offline settings. Nevertheless, effectively modeling diverse behavior policies in complex multi-agent environments remains a substantial challenge for offline MARL methods.
To address these challenges, we propose OffLight, an offline MARL framework specifically designed for handling heterogeneous behavior policies in TSC datasets. OffLight integrates importance sampling (IS) to correct distributional shifts and return-based prioritized sampling (RBPS) to highlight high-quality experiences. Additionally, OffLight uses a Gaussian mixture model variational graph autoencoder (GMM-VGAE) to represent diverse behavior policies accurately. Experiments across real-world urban traffic scenarios show that OffLight outperforms existing offline RL methods, demonstrating reduced average travel times and queue lengths. Our results highlight OffLight’s capability to reliably manage urban traffic networks without the risks associated with online training, making it a practical and scalable solution for real-world deployment.
II Background
II-A Multi-Agent Reinforcement Learning
We model the MARL problem as Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs) [16], formally defined by the tuple $\langle N,\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{N},\{\mathcal{O}^{i}\}_{i=1}^{N},P,R,H,\gamma\rangle$. Here, $N$ is the number of agents, $\mathcal{S}$ is the state space, $\mathcal{A}^{i}$ and $\mathcal{O}^{i}$ are the action and observation spaces for agent $i$, respectively, $P$ is the state transition probability, $R$ is the shared reward, $H$ is the finite horizon, and $\gamma\in[0,1)$ is the discount factor.
Each agent $i$ selects actions based on its local action-observation history $h^{i}_{t}=\{o^{i}_{0},a^{i}_{0},\dots,o^{i}_{t}\}$, resulting in joint action $\mathbf{a}_{t}=(a_{t}^{1},\dots,a_{t}^{N})$, transitioning the environment to state $s_{t+1}$ according to $P$, and yielding a shared reward $R$. The objective is to determine decentralized policies $\{\pi^{i}\}_{i=1}^{N}$, with each policy $\pi^{i}:H^{i}_{t}\rightarrow\mathcal{A}^{i}$ mapping histories to actions, to maximize expected cumulative reward [^1] In practice, we employ local rewards for each agent and adopt Centralized Training with Decentralized Execution (CTDE), as we found this approach converges faster than using a scalar global reward.:
@@ II-A Multi-Agent Reinforcement Learning
To optimize decentralized policies, Deep Q-Learning (DQNs) [17] is commonly employed, where each agent independently updates its Q-values based on local observations and rewards as follows: {dmath} Q^i(o^i_t,a^i_t) ←Q^i(o^i_t,a^i_t) + α(r^i_t + γmax_a’ Q^i(o^i_t+1,a’) - Q^i(o^i_t,a^i_t)).
However, independent learning may lead to suboptimal coordination. Recent frameworks, such as CoLight [11], PressLight [18], and AttendLight [19], address this by incorporating Graph Attention Networks (GATs) [20] for inter-agent communication. We leverage these successful frameworks as the basis for our approach.
II-B Offline Multi-Agent Reinforcement Learning
Offline Multi-Agent Reinforcement Learning (Offline MARL) learns policies from fixed datasets without additional environment interaction during training [3, 4].
Offline MARL presents several challenges: (1) Distributional Shift: The dataset may differ significantly from the policy’s actual environment interactions, leading to inaccurate value estimations and reduced performance [21]. (2) Extrapolation Error: Without interactive feedback, agents risk generalizing poorly to situations not covered by the dataset. (3) Limited Coverage: Historical data might not include optimal coordination strategies necessary for effective multi-agent policies.
Recent offline RL algorithms address these issues. Conservative Q-Learning (CQL) [12] introduces a penalty for overestimating Q-values of actions not in the dataset. Its objective is: {dmath} min_Q_i α( Eo_i ∼D, a_i ∼π(a_i|o_i) [Q(o_i,a_i)] - E(o_i,a_i) ∼D [Q(o_i, a_i)] ) + L_Bellman(Q_i).
TD3+BC [22] combines TD3 with a behavior cloning term to reduce policy deviation from the dataset, thus managing distributional shift: {dmath} min_π_θ_i -Eo_i ∼D [Qϕ(o_i, π_θ_i(o_i))] + λE(o_i,a_i) ∼D [|π_θ_i(o_i) - a_i|_2^2].
Both methods can be adapted to multi-agent systems using Centralized Training with Decentralized Execution (CTDE) [23, 24], enabling centralized training and decentralized agent execution for better scalability and coordination.
II-C Motivation for the OffLight Framework
Offline MARL faces unique challenges when datasets contain heterogeneous behavior policies, especially in domains like Traffic Signal Control (TSC). Different operational contexts or sources create diverse behaviors, increasing the likelihood of distributional mismatch. This diversity leads to unreliable Q-value estimates and degraded performance, as distributional correction methods become less effective [21]. Additionally, accurately estimating mixed behavior policies is difficult, complicating policy constraints and importance sampling [13].
To address heterogeneous behavior policies in offline MARL, we propose the OffLight framework. OffLight utilizes a behavior policy estimation technique (GMM-VGAE) to accurately model policy diversity, employs importance sampling to mitigate distributional shift, and integrates Reward-Based Prioritized Sampling (RBPS) for improved sample efficiency. RBPS prioritizes high-reward episodes, enhancing the focus on valuable training data. These strategies aim to improve policy robustness and performance specifically in Traffic Signal Control (TSC) scenarios. The OffLight framework details are introduced in the next section.
III OffLight Framework